Calculate the theoretical value of the each solution mixture (A, B, C, D) using ICE table or Henderson Hasselbalch equation aHow would you calculate the pH of solutions 1-4?
Answer #1

Calculation of ph Theoretically (1) 1Р, 25 сн, сгон гkim che certit, - CH₂ Cook (ug) - Ab to 1M (Initial) a - lag, Carl Ka -1- км Рии ки - и - - - в съпи Cencentration и и 1 LG *, І.fх “Г и 2 г /ух / 4 Сән) ,и» .24 х - 3 рон --Сон) • 2-22 pH = 14 – P(2) AM, arme Nhyou үлм и м 1, 2 • 27 Molar mans of N My CO . 14+ 4x + 58= r2. 14 + 4x14 1195 = 5.3,5g/mole polarity of NH, U